### v1

# Multiscale Modeling of Vacancy-Cluster Interactions and Solute Clustering Kinetics in Al-Mg-Zn Alloys

Abstract
During heat treatment of age-hardenable alloys, growing solute clusters can strongly trap excess vacancies, limiting their mobility and significantly influencing precipitation kinetics. However, quantitative understanding of vacancy-cluster interactions remains limited due to the inherently multi-time- and multi-length-scale nature of vacancy-mediated diffusion. In this work, we develop an integrated computational framework combining lattice kinetic Monte Carlo (KMC) simulations, an atomistic absorbing Markov chain model, and mesoscale cluster dynamics (CD) to investigate these interactions and predict their effects on long-time clustering in Al-Mg-Zn alloys. The Markov chain model yields vacancy escape times from solute clusters and identifies a two-stage behavior of the vacancy-cluster binding energy: a size-dependent regime for small clusters (fewer than $\sim$100 atoms), followed by a saturation stage governed by vacancy formation energy differences between the cluster and matrix. These binding energies are used to estimate residual mobile vacancy concentrations in the Al matrix after quenching, which serve as critical inputs to CD simulations to predict the long-time cluster evolution kinetics during natural aging. Our results quantitatively demonstrate that cooling rate substantially affects natural aging kinetics, offering predictive insights for tailoring heat treatments in engineering alloys.
CMC
canonical Monte Carlo
CNT
classical nucleation theory
CD
cluster dynamics
DFT
density functional theory
FCC
face-centered cubic
GP
Guinier-Preston
GB
grain boundary
KMC
kinetic Monte Carlo
MC
Monte Carlo
PEL
potential energy landscape
PF
phase-filed
SA
simulated annealing
SQS
special quasirandom structures
Introduction
[…]
These complex interactions between vacancies and growing solute clusters in multicomponent alloys remain relatively underexplored by computational methods due to their chemical complexity and the inherently multi-time- and multi-length-scale nature of the underlying diffusion processes. Atomistic methods, such as on-lattice kinetic Monte Carlo (KMC) simulations, can accurately model vacancy migration kinetics using local-environment-dependent migration barriers informed by first-principles calculations [14, 15, 16]. However, these methods face inherent limitations in both spatial and temporal scales. Spatial constraints arise from the need to explicitly track individual atomic and vacancy jumps [17, 18], while temporal limitations stem from frequent low-barrier events that drastically reduce simulation efficiency [19, 20, 21]. Furthermore, conventional on-lattice KMC approaches struggle to capture the dynamic evolution of mobile vacancy concentrations [8, 9, 10], which is essential for predicting long-term diffusion and phase transformation behavior. To address these challenges, computational methods, such as phase-filed (PF) simulations [22, 23, 24], are needed to capture the feedback between evolving vacancy concentrations and solute clustering on the mesoscale. However, these mesoscale models require inputs such as mobility functions and initial cluster configurations, which ideally should be derived from atomistic simulations based on actual alloy compositions and heat treatment histories. Yet, this connection remains difficult due to the aforementioned limitations of atomistic methods. Consequently, a significant knowledge gap persists in linking atomic-scale vacancy-mediated diffusion with mesoscale phase transformation modeling.
Bridging the knowledge gap in vacancy-mediated diffusion processes is particularly critical for 7000 series Al-Mg-Zn alloys. These alloys can achieve high strengths (up to approximately 700 MPa ultimate tensile strength) after appropriate heat treatments [25]. However, their broader application, such as in the automotive industry,is limited because the rapid formation of solute clusters during natural aging significantly reduces ductility and formability [26, 27, 28]. Our recent studies [29, 30] revealed that Mg-Zn-rich solute clusters form rapidly during the quenching process, rather than only after reaching room temperature. Lattice KMC simulations show that these early-formed clusters strongly trap vacancies, leading to a significant reduction in the mobile vacancy concentration and potentially slowing subsequent precipitation kinetics [30]. This phenomenon, often referred to as “vacancy prisons” [31, 32], corresponds to deep energy basins on the potential energy landscape (PEL) that govern vacancy migration [11]. Importantly, however, vacancies trapped in these clusters are not completely immobile. The vacancy migration barriers within Mg-Zn-rich clusters are generally lower than in the surrounding Al matrix, allowing relatively fast vacancy diffusion within the cluster region and enabling a finite probability for vacancies to escape back into the matrix [30].
Although these qualitative descriptions provide useful insight into vacancy trapping, the overall effects on long-time solute clustering kinetics are still unclear due to the limitations of atomistic scale methods described above. A quantitative understanding of vacancy-cluster interactions remains limited. Zurob and Seyedrezai [33] proposed a model wherein the vacancy trapping effect is proportional to the cluster size (i.e., the number of solute atoms), which has been validated for small clusters in Al alloys [34, 35]. This relates to the growing spatial extent of the energy basin. In contrast, Soisson et al. [10] introduced a different approach for Cu clusters in the Fe-Cu system, emphasizing that the binding energy is governed by the vacancy formation energy difference in the Cu precipitate and the Fe matrix, which dictates the energetic depth of the trapping basin. While these models capture different aspects of the PEL, their application to multicomponent systems, such as Al-Mg-Zn, is hindered by the combinatorial complexity of cluster configurations and the complex interactions among different solute species. As a result, accurately modeling vacancy-cluster interactions in such systems remains a significant challenge, and conventional approaches (e.g., simple binding energy estimates) may not sufficiently account for this complexity [36, 11].
Fortunately, similar challenges involving low-energy basins have been extensively studied in the context of accelerating KMC simulations [37, 38, 39, 40, 41, 42, 43]. Conventional KMC approaches often become inefficient when systems are repeatedly trapped in states separated by small energy barriers, leading to prolonged residence in low-energy basins. To address this issue, several advanced methods have been developed. In particular, Chatterjee et al. [19], Puchala et al. [20], and Fichthorn et al. [21] introduced absorbing Markov chain frameworks that accelerate KMC simulations by efficiently treating transitions within such basins. Building upon these strategies, we apply an absorbing Markov chain model [20, 21] to systematically quantify the low-barrier, low-energy basins associated with Mg-Zn-rich solute clusters in Al-Mg-Zn alloys. In this approach, solute clusters are modeled as bidirected graphs, where nodes represent possible vacancy positions and edges denote transitions between nearest-neighbor lattice sites. By analyzing the transition probabilities and absorbing state behavior, we compute vacancy escape times from clusters and validate our results against direct KMC simulations. Additionally, we evaluate how effective vacancy binding energies vary with cluster size and assess their influence on the concentration of residual mobile vacancies in the Al matrix, a critical factor governing long-time diffusion kinetics after quenching. Our Markov chain model reveals a two-stage behavior in vacancy-cluster binding energies: an initial size-dependent regime for small clusters, consistent with the model proposed by Zurob and Seyedrezai [33], followed by a saturation regime governed by the vacancy formation energy difference between the cluster and matrix, in line with the findings of Soisson et al. [10]. Further details are provided in the Results section.
Finally, by leveraging the effective vacancy binding energies as functions of cluster size, which is quantified through the integrated KMC and absorbing Markov chain framework described above, we obtain essential input parameters for mesoscale simulations to capture long-time diffusion kinetics and associated diffusional phase transformations. cluster dynamics (CD) simulations provide a mesoscale modeling approach to describe the temporal evolution of atomic clusters (e.g., solute-vacancy clusters, precipitates, or defect aggregates) using population-based rate equations, without explicitly tracking individual atoms [44, 45, 46, 47, 48]. These rate equations incorporate the vacancy-cluster binding energy functions derived from our atomistic framework. Accordingly, we construct CD models to investigate long-time clustering kinetics, allowing mobile vacancy concentrations to evolve dynamically with the solute cluster configurations. The predicted results correspond well to experimentally relevant time and length scales. Our findings quantitatively demonstrate that cooling rate strongly influences natural aging kinetics, offering predictive insights for optimizing heat treatments in engineering alloys. Additionally, the predicted cluster distributions provide critical initial conditions for solute nucleation and early-stage growth in PF simulations of precipitation kinetics [22, 23, 24].
Results
Kinetic Monte Carlo simulations during quenching
Figure 1: Time evolution of the average energy change per atom $E$ of two typical KMC quenching simulations[30]. The curve above is at a constant temperature of 300 K from the first step, representing an infinitely fast (instantaneous) cooling simulation from a solid solution. The curve below is from a water-quenching simulation. The change in the simulation temperature is denoted by the colors defined in the right color bar. Snapshots illustrate the initial and final structures from each simulation. The length of the cubic simulation supercells in the simulation snapshots is 12 nanometers.
[…]
Another significant difference between the instantaneous cooling and water quenching cases in the above KMC simulations lies in the solute clustering kinetics at low temperatures ($\sim$300 K). In the instantaneous cooling case, the system’s energy continues to decrease at 300 K, indicating ongoing, albeit slow, vacancy-mediated solute diffusion. In contrast, energy evolution ceases entirely in the water quenching case, suggesting that vacancies become effectively trapped by pre-formed solute clusters.
Energetic analysis of vacancy-cluster interactions
To systematically study the mechanism behind vacancy trapping in Al-Mg-Zn alloys, we first aimed to determine the structures and energetics of the clusters. Although density functional theory (DFT) calculations offer reliable energetic estimations, identifying the most stable structures of clusters with the given $n_{\text{Mg}}$ and $n_{\text{Zn}}$ in Al-Mg-Zn multicomponent alloys remains challenging because the possible configurations of solute clusters explode exponentially with the increase of $n_{\text{Mg}}$ and $n_{\text{Zn}}$ due to the change of atomistic structures. Hence, we employed surrogate models trained on first-principles data (see Supplementary Note 1) and applied a simulated annealing (SA) heuristic [49] adapted from the Metropolis-Hastings algorithm [50]. This method efficiently explores configuration space while minimizing the average formation enthalpy per solute atom, denoted as $\bar{H}_{\text{form}}$ (see Supplementary Note 2).
Figure 2: Analyses of cluster local energetics obtained from SA. (a)-(c) An example cluster presenting the inner structures. The cluster is colored to show the spatial distribution of element type, the local vacancy binding energy of each site $E^{\text{site}}_{\text{bind}}$, and the local average vacancy migration barrier of each site $E^{\text{site}}_{\text{mig}}$. (d) Kernel density estimation plot for local vacancy migration barrier of each bond $E^{\text{bond}}_{\text{mig}}$ of this cluster based on their migration type. Here, NN shell denotes the first nearest neighbor shell of the cluster. The yellow region marks the vacancy migration event inside clusters. Light blue denotes the event vacancy jumping from the surface of the cluster to the nearest neighbor shell. Dark blue represents the event vacancy jumping out of the nearest neighbor shell. Dashed lines denote the averaged value for each type, and the black dotted line denotes the vacancy migration barrier in pure Al, $E_{\text{mig}}^{\text{Al}}$, of 0.58 eV. (e) Conceptual energy landscape of a ‘vacancy prison’ generated by a solute cluster. The dashed line connects local minima, while the curved line connects both minima and transition states.
After the structures of the clusters were obtained, we conducted a comprehensive analysis. Specifically, we examined the local vacancy binding energy at each lattice site, $E^{\text{site}}_{\text{bind}}$, and the local migration barriers associated with each atomic bond, $E^{\text{bond}}_{\text{mig}}$. Figs. 2a-c depict an example cluster extracted from the SA simulations, sliced to show its internal structure. Fig. 2a highlights the atomic arrangement of Al (blue), Mg (green), and Zn (orange). This cluster exhibits a structure similar to $L1_{2}$, but a Zn/Mg ratio of approximately 2 due to its high surface-area-to-volume ratio. Fig. 2b presents the spatial distribution of the local binding energy of each site, $E_{\text{bind}}^{\text{site}}$, calculated from the following expression:
$E_{\text{bind}}^{\text{site}}=\left[E\left(\text{Vac}\right)-E\left({\text{Al}}_{N-1}\text{Vac}\right)\right]-\left[E\left(\text{X}\right)-E\left({\text{Al}}_{N-1}\text{X}\right)\right],$ (1)
where $E(\text{Vac})$ denotes the energy of the supercell containing a vacancy at a particular site in the cluster, and $E(\text{X})$ represents the energy of the same supercell with the atom X to fill up the previous local vacancy site. Atom X is the element originally present at that location, which can be Al, Mg, or Zn. Similarly, $E({\text{Al}}_{N-1}\text{Vac})$ and $E({\text{Al}}_{N-1}\text{X})$ correspond to the energies of a larger, pure Al supercell with a vacancy and with atom X occupying the vacancy site, respectively. This equation compares the formation energy difference of a vacancy-free cluster to the same cluster but containing a vacancy at a particular site. The value of $E_{\text{bind}}^{\text{site}}$ quantifies how strongly a vacancy binds at a particular site relative to the solute that occupies the position. Specifically, in a dilute Al matrix, $E_{\text{bind}}^{\text{site}}=0$, indicating no additional binding for the vacancy. In contrast, within a cluster, $E_{\text{bind}}^{\text{site}}$ typically exhibits a negative value, suggesting that vacancies are more likely to be bound or trapped within the cluster compared to the Al matrix. Fig. 2c similarly presents the spatial distribution of the local average migration barriers at each site, $E_{\text{mig}}^{\text{site}}$. This value represents the arithmetic average of the 12 vacancy migration barriers, $E_{\text{mig}}^{\text{bond}}$, associated with bonds connecting to its 12 first-nearest neighbors, assuming the site is initially occupied by a vacancy. The results highlight numerous sites within the cluster that exhibit low average migration barriers, with values around $\sim 0.3$ eV. Vacancies located at these sites have notably lower migration barriers compared to those at sites near the cluster shell, which present higher values similar to the vacancy migration barrier in pure Al, $E_{\text{mig}}^{\text{Al}}$, of $0.58$ eV.
To further investigate the variation in migration barriers, we classify the migration barriers for each bond, $E_{\text{mig}}^{\text{bond}}$, into three distinct groups, that is, migrations within clusters (yellow), transitions from the cluster surface to the nearest neighbor shell (light blue), and jumps beyond the nearest neighbor shell (dark blue), as depicted in Fig. 2d. The distributions of $E_{\text{mig}}^{\text{bond}}$ for each group are plotted. Consistent with observations in Fig. 2c, migration barriers are typically lower within the cluster. In contrast, vacancies jumping out of the nearest neighbor shell exhibit barriers primarily in the range of 0.54 to 0.62 eV, closely matching the vacancy migration barrier in pure Al. Ultimately, Fig. 2 illustrates that the cluster’s energetics create a low-energy basin in the PEL, as depicted in Fig. 2e. The basin presents mainly a short-range effect, and the cluster-matrix boundary can be interpreted as defining the edge of this energy basin, at which $E_{\text{bind}}^{\text{site}}\approx 0$ and $E_{\text{mig}}^{\text{bond}}\approx E_{\text{mig}}^{\text{Al}}$.
Absorbing Markov chain model for vacancy escape from clusters
Figure 3: Workflow for constructing an absorbing Markov chain model to analyze vacancy-cluster interactions and predict vacancy escape times. (a) An example solute cluster showing atomic arrangements and element distributions, with colors indicating Al (blue), Mg (green), and Zn (orange). (b) A graph representation of the cluster, including an additional layer of neighboring atoms. Nodes correspond to atoms/vacancies at lattice sites, and edges depict bonds between nearest neighbors. (c) An absorbing Markov chain (bidirected graph) derived from the directed graph, with nodes colored by local binding energies $E_{\text{bind}}^{\text{site}}$ and edges shaded based on migration barriers between connected nodes (sites) $E_{\text{mig}}^{\text{bond}}$. The outermost layer represents absorbing states where vacancies are considered escaped.
Therefore, we construct a vacancy escape model based on an absorbing Markov chain framework (see Methods section for concepts), based on the energetics obtained from the surrogate model (Figs. 2b-c). The workflow used to build the absorbing Markov chain model, predict vacancy escape times, and analyze vacancy-cluster interactions is illustrated in Fig. 3. As depicted in Fig. 3a, it starts with solute clusters. We obtained cluster structures directly from SA, although clusters can similarly be derived from KMC or other atomistic simulation methods. Then, the atomic structure is transformed into a graph representation (Fig. 3b). Each atom is represented as a node, and each atomic bond as an edge. To accurately describe the event vacancy jumping out of the nearest neighbor shell (Fig. 2d), an additional layer of neighboring atoms surrounding the cluster is included. Next, the graph is parameterized to become a bidirected graph, as shown in Fig. 3c. This transformation introduces directionality into the graph, labeling the information of vacancy migration paths. The outermost layer in this bidirected graph represents absorbing states, which are viewed as states in the Al matrix, where vacancies are considered to have escaped from the cluster region. The local binding energy, $E_{\text{bind}}^{\text{site}}$, is calculated for each node to characterize the thermodynamic stability of vacancies at specific lattice sites. Similarly, vacancy migration barriers between connected nodes, $E_{\text{mig}}^{\text{bond}}$, are also computed to parameterize transition probabilities between different sites.
Nodes (lattice sites) in the cluster and the nearest neighbor shell (lattice sites as the nearest neighbors of any sites in the cluster) are treated as transient states, whereas nodes (lattice sites) in the outermost layer (off the nearest neighbor shell) act as absorbing states. Transient states represent sites that vacancies can temporarily occupy and subsequently migrate away from, whereas absorbing states denote sites that are considered as the vacancy having escaped the cluster. Fig. 3c presents the same information as Figs. 2b-c: interior nodes generally exhibit lower (more negative) binding energies, signifying stronger vacancy binding, while bonds inside the cluster typically show lower migration barriers, highlighting easier vacancy migration within the cluster.
Finally, the absorbing Markov chain model is constructed based on the graph representation of Fig. 3c. The mean exit time that a vacancy escapes from the cluster, $t_{\text{esc}}$, can be determined using the formula [21]:
$t_{\text{esc}}=\vec{p_{0}}^{T}(\boldsymbol{I}-\boldsymbol{T})^{-1}\vec{\tau},$ (2)
Here $\boldsymbol{T}$ is the transient matrix containing transition probabilities among all transient states, which is a square matrix of size $n_{\text{tr}}\times n_{\text{tr}}$, where $n_{\text{tr}}$ is the number of transient states and determined by sites in the cluster and the nearest neighbors. $\boldsymbol{I}$ is an identity matrix that has the same size as $\boldsymbol{T}$. $\vec{p_{0}}$ represents the initial probabilities of transient states with size $n_{\text{tr}}\times 1$, and $\vec{\tau}$ is a vector of one-step escape times for each site, which has the same size as $\vec{p_{0}}$.
Specifically, elements $T_{ij}$ from the matrix $\boldsymbol{T}$ are derived from the migration barriers $E^{\text{bond}}_{\text{mig}}$ based on an Arrhenius relationship:
$T_{ij}=\frac{\exp\left(-\frac{E^{\text{bond}}_{\text{mig},ij}}{k_{B}T}\right)}{\sum_{j=1}^{z}\exp\left(-\frac{E^{\text{bond}}_{\text{mig},ij}}{k_{B}T}\right)},$ (3)
where $z$ is the coordination number ($z=12$ for an face-centered cubic (FCC) lattice), $E^{\text{bond}}_{\text{mig},ij}$ is the migration barrier from node (lattice site) $i$ to node (lattice site) $j$, corresponding to one of the 12 vacancy migration barriers from a lattice site to the nearest neighbor sites used to calculate previously defined $E_{\text{mig}}^{\text{bond}}$. To obtain the values of elements $p_{i}$ from $\vec{p_{0}}$, an equilibration assumption has been employed [20, 21, 42]. It assumes that each transient state will be visited numerous times since the energy basin cannot be escaped easily, eventually reaching a local equilibrium within the basin. Once this local equilibrium is established, the probability of occupying any given transient state $i$ follows a Boltzmann distribution:
$p_{i}=\frac{\exp\left({-\frac{E^{\text{site}}_{\text{bind},i}}{k_{B}T}}\right)}{\sum_{j}\exp\left({-\frac{E^{\text{site}}_{\text{bind},i}}{k_{B}T}}\right)},$ (4)
where $E^{\text{site}}_{\text{bind},i}$ is the energy of vacancy on site $i$. The elements, $\tau_{i}$, in the vector $\vec{\tau}$ represent the expected escape one-step time for the particular site. For each site $i$, $\tau_{i}$ accounts for all possible transitions to neighboring states. Migration barriers $E_{a{ij}}$, and attempt frequencies $\nu$ are incorporated to estimate escape times from each transient state.
$\tau_{i}=\left(\nu\sum_{j}{\exp\left(-\frac{E^{\text{bond}}_{\text{mig},ij}}{k_{B}T}\right)}\right)^{-1}.$ (5)
Here $\nu$ is set at $10^{13}$ Hz, as a typical atomic vibration frequency value in a crystal.
Figure 4: KMC simulations for the verification of the absorbing Markov chain model. (a) The strategy for applying the KMC algorithm to determine the escaped state of a vacancy from a solute cluster. Open red circles denote vacancies. Blue, green, and orange circles represent Al, Mg, and Zn atoms, respectively. White stars mark the first nearest neighbor shell of the cluster. (b) The results comparison of the escape time $t_{\text{esc}}$ defined in Eq. 2 from the absorbing Markov chain model and from KMC verification for solute clusters generated by SA with different cluster sizes $N^{\text{clu}}$.
These energetics applied in the absorbing Markov chain model, $E_{\text{bind}}^{\text{site}}$ and $E_{\text{mig}}^{\text{bond}}$, provide sufficient information to construct an accurate PEL for single-component clusters, such as Cu clusters in Fe-Cu systems, or Cu Guinier-Preston (GP) zones in Al-Cu alloys. In such systems, vacancy migration inside clusters does not induce structural rearrangements due to the indistinguishable nature of solute atoms. However, for binary clusters in multicomponent systems, such as Mg-Zn clusters in Al-Mg-Zn alloys, we recognize that local atomic rearrangements by vacancy migration may complicate the actual PEL. Despite these structural changes, we make a simplified-PEL assumption about the PEL. Specifically, we assume that the PEL experienced by a migrating vacancy can be approximated by the static ground-state structures of the cluster (obtained by SA). This simplification is reasonable at lower temperatures, where thermodynamics favors highly ordered, low-energy structures. By making this assumption, we effectively reduce the complexity of the PEL, enabling a computationally efficient absorbing Markov chain model to describe vacancy dynamics inside the clusters. This simplification maintains analytical tractability and computational efficiency, effectively capturing essential vacancy behavior without explicitly modeling structural rearrangements.
To validate the absorbing Markov chain model and test the assumption of a simplified PEL, we applied KMC simulations. KMC and Markov chains share the same memorylessness property, which assumes that the next state depends only on the current state and not on past states. KMC calculates transition rates and selects events probabilistically, updating the system state and time step by step. Unlike the Markov model, KMC inherently accounts for structural rearrangements and complex PELs by simulating vacancy migration.
Fig. 4 presents the verification for the clusters generated and optimized from the SA method (see Supplementary Note 3 for the clusters extracted from KMC simulations). Each cluster contains one vacancy and maintains a fixed Zn/Mg ratio of 2. Then, we performed KMC simulations on these clusters, focusing on determining the real escape times required for a vacancy to leave the cluster. In KMC simulations, the vacancies are tracked as they migrate through the cluster until they reach the surface, where they are considered to have escaped. This stopping criterion, illustrated in Fig. 4a, mirrors the definition used in the absorbing Markov chain model, providing a consistent basis for comparison between the two approaches. The KMC simulations were conducted at an elevated temperature of 450 K. This higher temperature facilitates fast transitions, enabling vacancies to overcome energy barriers and escape the energy basin more efficiently compared to simulations performed at room temperature. This higher temperature ensures that statistically meaningful results can be obtained within a feasible computational time. Fig. 4b compares the escape times predicted by the absorbing Markov chain model (Eq. 2) with those obtained from the KMC simulations. Here, results from absorbing Markov chain and KMC validation exhibit strong agreement, demonstrating that the simplified PEL approximation employed in the absorbing Markov model effectively captures the essential features of vacancy dynamics.
Vacancy trapping capacity of solute clusters
Figure 5: Analysis for vacancy trapping ability of solute clusters using the Markov chain model. (a) Estimated vacancy escape time $t_{\text{esc}}$ at 300 K of solute clusters in terms of cluster size $N^{\text{clu}}$. (b) A conceptual figure to explain the singularity approximation for an energy well generated by a solute cluster. (c) Effective vacancy binding energy of a cluster $\tilde{E}^{\text{clu}}_{\text{bind}}$ defined in Eq. 7 in terms of cluster size $N^{\text{clu}}$.
Accordingly, the estimated vacancy escape times at 300 K, denoted as $t_{\text{esc}}$ and calculated using Eq. 2, are presented in Fig. 5a. The general trend of $t_{\text{esc}}$ with cluster size broadly exhibits two distinct stages, each reflecting a different vacancy trapping characteristic. In the initial stage, as cluster size increases, vacancy escape time increases sharply on a logarithmic scale, from approximately $10^{-3}$ seconds to $10^{1}$ seconds. This rapid growth indicates that smaller clusters significantly enhance their vacancy trapping ability with growth. This trend aligns with the observation that the energy basin depth (i.e. the lowest site energy within the cluster) increases linearly with the size for smaller clusters [33].
However, as cluster size exceeds roughly 100 atoms, this trend transitions to a second stage, where further increases in cluster size lead to only marginal changes in $t_{\text{esc}}$. This behavior indicates that the vacancy trapping becomes relatively insensitive to size for large clusters. The underlying reason is that the depth of the PEL basin saturates, converging to the vacancy formation free energy difference between the Al matrix and the solute-rich secondary phase [8, 10, 11]. Beyond this point, continued growth increases only the width rather than the depth of the energy basin. Although the width effect, associated with additional random walk steps that a vacancy must take to reach the cluster boundary, is typically less dominant than the depth effect due to relatively low internal migration barriers, it is not negligible. As the number of internal steps increases, we still observe that the vacancy escape time continues to grow.
Now, our absorbing Markov chain model captures both the depth and the width of the energy basin, incorporating the combined effects of both thermodynamic binding energies and kinetic migration barriers within the cluster. With the vacancy escape times (Fig. 5a), we can more comprehensively characterize the vacancy trapping mechanisms in terms of energetics. We adopt a singularity approximation of the PEL, illustrated in Fig. 5b, where the cluster is simplified as a single site with strong vacancy binding instead of representing the cluster as a complex energy basin of the PEL. This simplification allows us to quantify the trapping ability of the cluster through a single effective binding energy value.
We define the effective cluster energy barrier, $\tilde{E}^{\text{clu}}_{\text{barrier}}$, which governs the escape time via the Arrhenius relationship:
$t_{\text{esc}}=\nu^{-1}\exp\left(\frac{\tilde{E}^{\text{clu}}_{\text{barrier}}}{k_{B}T}\right).$ (6)
By inverting Eq. 6 and using $E_{\text{mig}}=E_{\text{mig}}^{\text{Al}}=0.58$ eV as a typical vacancy migration barrier outside of the basin, we define the effective vacancy binding energy, $\tilde{E}^{\text{clu}}_{\text{bind}}$, as:
$\tilde{E}^{\text{clu}}_{\text{bind}}=E_{\text{mig}}-\tilde{E}^{\text{clu}}_{\text{barrier}}=E_{\text{mig}}^{\text{Al}}-k_{B}T\ln(\nu t_{\text{esc}}).$ (7)
As indicated by Fig. 5b, $\tilde{E}^{\text{clu}}_{\text{barrier}}$ can be larger than $E_{\text{mig}}$ so that a negative value of $\tilde{E}^{\text{clu}}_{\text{bind}}$ indicate that a vacancy is effectively trapped by a solute cluster. Fig. 5c plots $\tilde{E}^{\text{clu}}_{\text{bind}}$ as a function of the cluster size, with dashed lines indicating numerical fits. These results quantitatively confirm that for large clusters, the $\tilde{E}^{\text{clu}}_{\text{bind}}$ increases rapidly and approaches a size-independent limit due to the saturation of binding strength within the interior. Nevertheless, even in this saturation stage, the effective binding energy continues to capture the impact of the random walk behavior of the vacancies within the cluster.
Impacts on residual vacancy concentrations for diffusion
In multicomponent alloys, the vacancy concentration is influenced not only by heat treatment conditions but also by the chemical composition and spatial distribution of solute atoms. These solute atoms can aggregate to form clusters that effectively bind vacancies, thereby altering the concentration of mobile vacancies in the lattice. The effective binding energy to a given solute cluster, denoted as $\tilde{E}^{\text{clu}}_{\text{bind}}$, derived from the absorbing Markov chain model, offers a rigorous and quantitative framework for evaluating the impact of various types and sizes of solute clusters on excess vacancy concentrations. Yang et al. [51] reviewed prior studies [8, 9, 10, 52, 53] and summarized the equilibrium vacancy concentration as a function of temperature. In this framework, the total number of vacancies is categorized into three distinct groups:
$c^{\text{vac}}_{\text{eq}}\left(T\right)=c^{\text{vac,X}}_{\text{eq}}+c^{\text{vac,clu}}_{\text{eq}}+c^{\text{vac,res}}_{\text{eq}}.$ (8)
The first group, $c^{\text{vac,X}}_{\text{eq}}$, denotes the atomic fraction of vacancies trapped by individual solute atom X. The reference vacancy concentration in pure Al is represented by $c^{\text{vac,Al}}_{\text{eq}}$. Both this quantity and the concentration of isolated solute atoms, $c^{\text{X}}_{1}$, are expressed as atomic fractions. The parameter $z$ represents the coordination number, i.e., the number of nearest-neighbor sites surrounding a solute atom. The term $E_{\text{bind}}^{\text{X}}$ denotes the vacancy binding energy associated with solute atom X.
$c^{\text{vac,X}}_{\text{eq}}=c^{\text{vac,Al}}_{\text{eq}}\left(T\right)\sum_{\text{X}}\left[zc^{\text{X}}_{1}\exp\left(\frac{E_{\text{bind}}^{{\text{X}}}}{k_{B}T}\right)\right],$ (9)
The second group, $c^{\text{vac,clu}}_{\text{eq}}$, is the fraction of vacancies trapped by clusters. The sizes of these clusters are expressed as $N^{\text{clu}}$, which is the number of lattice sites occupied by this cluster. The concentrations of these clusters, $c^{\text{clu}}$, are expressed as an atomic fraction (number of clusters in the whole simulation supercell divided by the number of lattice sites in this supercell). The term $z^{\text{clu}}$ represents the number of neighboring sites which are direct nearest neighbors to any atom belonging to the cluster. These sites provide additional sites for trapping vacancies on the cluster’s surface. $\tilde{E}_{\text{bind}}^{\text{clu}}$ is the effective vacancy binding energy of clusters that can be provided by Eq. 7.
$c^{\text{vac,clu}}_{\text{eq}}=c^{\text{vac,Al}}_{\text{eq}}\left(T\right)\sum_{\text{clu}}\left[\left(z^{\text{clu}}+N^{\text{clu}}\right)c^{\text{clu}}\exp\left(\frac{\tilde{E}_{\text{bind}}^{\text{clu}}}{k_{B}T}\right)\right],$ (10)
The last group, $c^{\text{vac,res}}_{\text{eq}}$, represents the residual vacancies that remain freely distributed in the Al matrix.
$c^{\text{vac,res}}_{\text{eq}}=c^{\text{vac,Al}}_{\text{eq}}\left(T\right)\left[1-\sum_{\text{X}}{\left(z+1\right)c^{\text{X}}}-\sum_{\text{clu}}\left(z^{\text{clu}}+N^{\text{clu}}\right)c^{\text{clu}}\right],$ (11)
Although Eq. 11 offers a refined estimate of the equilibrium residual vacancy concentration, the vacancy concentration during quenching remains far from equilibrium due to rapid kinetic processes. Recent computational and experimental studies [4, 5] have demonstrated that vacancy supersaturation can persist for several hours to days ($\sim 10^{5}$-$10^{6}$ seconds) in regions far from GB. Therefore, we focus on the interactions between vacancies and both solute atoms and solute clusters, while assuming that the total vacancy concentration remains constant during quenching in large-grained regions. The expression for the dynamic vacancy concentration is given by:
$c^{\text{vac,res}}_{\text{dyn}}=c^{\text{vac}}_{\text{eq}}\left(800K\right)\frac{\sum_{\text{Al}}\exp\left({-\frac{E_{\text{bind}}^{\text{site}}}{k_{B}T}}\right)}{\sum_{\text{all}}\exp\left({-\frac{E_{\text{bind}}^{\text{site}}}{k_{B}T}}\right)},$ (12)
where the numerator and denominator represent the total probability of finding vacancies in all lattice sites of the Al matrix and in all lattice sites of the entire alloy, respectively, where $E_{\text{bind}}^{\text{site}}$ is the vacancy binding energy of each lattice site. It is important to note that the local binding energy of sites in the pure Al matrix is defined to be zero. By substituting the effective cluster binding energy, $\tilde{E}_{\text{bind}}^{\text{clu}}$, for $E_{\text{bind}}^{\text{site}}$ for each site belong to clusters, Eq. 12 can be reformulated as:
$c^{\text{vac,res}}_{\text{dyn}}=c^{\text{vac}}_{\text{eq}}\left(800K\right)\frac{1-\sum_{\text{clu}}\left(z^{\text{clu}}+N^{\text{clu}}\right)c^{\text{clu}}}{1-\sum_{\text{clu}}\left(z^{\text{clu}}+N^{\text{clu}}\right)c^{\text{clu}}+\sum_{\text{clu}}\left[\left(z^{\text{clu}}+N^{\text{clu}}\right)c^{\text{clu}}\exp\left(\frac{\tilde{E}_{\text{bind}}^{\text{clu}}}{k_{B}T}\right)\right]}.$ (13)
The term $c^{\text{vac,res}}_{\text{dyn}}$ represents the concentration of mobile vacancies remaining in the Al matrix that are available to facilitate solute diffusion. This model enables more accurate prediction of vacancy concentrations, particularly in systems containing large solute clusters.
Fig. 6a displays the temperature evolution profiles under three different cooling conditions: instantaneous quenching, fast cooling, and slow cooling (see Supplementary Note 4 for cluster snapshots). Fig. 6b presents the corresponding evolution of residual vacancy concentrations, including both the equilibrium concentration $c^{\text{vac,res}}_{\text{eq}}$ (from Eq. 11) and the dynamic concentration $c^{\text{vac,res}}_{\text{dyn}}$ (from Eq. 13) under these three conditions. In all cases, $c^{\text{vac,res}}_{\text{dyn}}$ remains higher than $c^{\text{vac,res}}_{\text{eq}}$ because equilibrium has not yet been reached. For the instantaneous quenching case, solute clusters cannot grow rapidly, leading to fewer large clusters capable of effectively trapping vacancies. As a result, a higher $c^{\text{vac,res}}_{\text{dyn}}$ is observed in the Al matrix. Over time, as clustering progresses and larger clusters form, $c^{\text{vac,res}}_{\text{dyn}}$ gradually decreases. In contrast, under more realistic cooling conditions (fast and slow cooling), stronger vacancy-cluster interactions are observed due to the formation of larger clusters during the cooling process. These clusters absorb more vacancies through strong binding interactions, resulting in lower $c^{\text{vac,res}}_{\text{dyn}}$ at a given time. Overall, $c^{\text{vac,res}}_{\text{dyn}}$ tends to be higher for faster cooling rates than for slower ones, as faster cooling suppresses the formation of additional large clusters that trap vacancies more effectively. Interestingly, this trend is opposite to that observed for $c^{\text{vac,res}}_{\text{eq}}$, where faster cooling corresponds to a lower temperature at a given time, thereby yielding lower equilibrium vacancy concentrations. This opposite trend can be rationalized by noting that faster cooling rates lead to a larger deviation between the dynamic and equilibrium vacancy concentrations, i.e., a greater gap between $c^{\text{vac,res}}_{\text{dyn}}$ and $c^{\text{vac,clu}}_{\text{eq}}$, due to the system being driven further from equilibrium by rapid kinetics.
Figure 6: Evolution of residual vacancy concentration during cooling, estimated from KMC simulations under three different cooling rates. (a) Temperature profiles corresponding to the three cooling conditions used in the KMC simulations. (b) Comparison of equilibrium residual vacancy concentration ($c^{\text{vac,res}}_{\text{eq}}$ defined in Eq. 11, dashed lines) and residual vacancy concentration ($c^{\text{vac,res}}_{\text{dyn}}$ defined in Eq. 13, solid lines) over time for instantaneous, fast, and slow cooling scenarios.
Cluster dynamics simulations for long-time clustering kinetics
Predictions of $c^{\text{vac,res}}_{\text{dyn}}$, the concentration of mobile vacancies available for solute diffusion in the Al matrix, provide a solid foundation for studying the nucleation and growth kinetics of solute clusters in Al alloys. The diffusivity of solute species, $D^{\text{X}}$, is directly influenced by vacancy concentration and migration barriers, and follows the relationship $D^{\text{X}}\propto c^{\text{vac,res}}_{\text{dyn}}\exp\left(-{E_{\text{mig}}^{\text{X}}}/{k_{B}T}\right)$. This diffusivity governs both the nucleation rate, $dN/dt\propto D^{\text{X}}\exp(-G^{*}/{k_{B}T})$, and the growth rate, $dR/dt\propto D^{\text{X}}/R$, which are critical for describing precipitation kinetics. The KMC results shown in Fig. 1 demonstrate that traditional KMC simulations encounter challenges in efficiently simulating cluster growth beyond the early stages. These limitations arise due to strong vacancy-cluster interactions and the finite number of vacancies in a typical simulation supercell. To overcome these constraints and to study clustering behavior over longer timescales relevant to natural aging, CD simulations were employed. By incorporating the accurately determined vacancy supersaturation obtained from the absorbing Markov chain model, the CD framework (see Methods section for details) bridges atomistic-scale vacancy interactions with macroscopic precipitation processes, enabling robust modeling of cluster size distribution evolution over extended time periods.
Figure 7: Comparison between CD and KMC simulations under infinitely fast (instantaneous) cooling conditions. (a) Time evolution of system energy per atom. Results are shown for KMC (blue dashed, light blue shade represents the standard error from three individual KMC runs), CD without vacancy-cluster interaction corrections (NVC, green), and CD with vacancy-cluster interaction corrections (VC, orange). (b)-(d) The distributions of cluster concentrations $c^{\text{clu}}$ as a function of cluster size $N^{\text{clu}}$ at the time of (b) 1 s, (c) 30 s, and (d) 880 s after the starting of the simulations.
To validate the predictive capability of the CD model, particularly its incorporation of vacancy-cluster interaction corrections in solute diffusivity, we directly compare the CD results with KMC simulations under infinitely fast (instantaneous) cooling conditions. Fig. 7a presents the time evolution of the system energy per atom, showing three datasets: KMC simulations, CD without vacancy-cluster interaction corrections (CD-NVC), and CD with corrections (CD-VC). The CD-VC model includes the effects of dynamic residual vacancy concentration, $c^{\text{vac,res}}_{\text{dyn}}$, which captures vacancy supersaturation and its influence on solute diffusivity $D^{\text{X}}$. Since solute diffusivity directly governs condensation and evaporation rates in the CD framework, accurately incorporating $c^{\text{vac,res}}_{\text{dyn}}$ allows CD-VC simulations to closely reproduce the energy evolution obtained from KMC over time. In addition, Figs. 7b-d compare the distributions of cluster concentrations, $c^{\text{clu}}$, as a function of cluster size, $N^{\text{clu}}$, obtained from CD-VC predictions and KMC simulations at different time points: 1 s, 30 s, and 880 s. These comparisons demonstrate strong agreement between the CD-VC and KMC models in predicting the temporal evolution of cluster size distributions across a wide range of times and temperatures.
Figure 8: Long-time CD simulation results of natural aging solute clustering for alloys obtained from different cooling conditions. (a) Residual vacancy concentration evolution during natural aging from different cooling conditions. (b)-(c) The distributions of cluster concentrations $c^{\text{clu}}$ as a function of cluster size $N^{\text{clu}}$ for various natural aging time (as-quenched(10 seconds), 1 hour, 6 hours, and 12 hours) for the alloys obtained from fast cooling (b) and slow cooling (c) conditions. The temperature profiles for the fast and slow cooling cases are described in Fig. 6(a).
Finally, the combination of KMC simulations and the CD model, informed by the calculated dynamic vacancy concentrations $c^{\text{vac,res}}_{\text{dyn}}$ (which directly influence solute diffusivity), is employed to simulate the long-time natural aging process following various cooling conditions, beginning from the as-quenched state. In this hybrid approach, KMC simulations are first used to model the quenching process until the temperature reaches approximately 300 K (corresponding to 10 seconds in our cases). At this point, the system energy evolution plateaus as shown in Fig. 1, indicating that most vacancies in the supercell are effectively trapped by solute clusters, and the system reaches a metastable state. Subsequently, the CD model is applied to simulate the long-term behavior of solute cluster evolution during natural aging. Fig. 8a presents the evolution of residual vacancy concentrations, $c^{\text{vac,res}}_{\text{dyn}}$, during natural aging at 300 K for both fast and slow cooling scenarios. These scenarios correspond to the two temperature profiles shown in Fig. 6a. The initial value of $c^{\text{vac,res}}_{\text{dyn}}$ at the onset of aging (i.e., at time $t=10$ s, marking the end of quenching) strongly depends on the cooling rate. Faster cooling limits the time available for solute clusters to form and trap vacancies, resulting in higher initial values of $c^{\text{vac,res}}_{\text{dyn}}$, as shown in Fig. 6b.
During the subsequent natural aging process, the dynamic residual vacancy concentration, $c^{\text{vac,res}}_{\text{dyn}}$, continues to decrease over time for both cooling scenarios. This decline reflects the gradual annihilation of excess vacancies at various sinks, particularly growing solute clusters. Although dislocations and GB are also known to act as vacancy sinks, they typically become significant only at much longer timescales (on the order of $10^{4}$-$10^{5}$ s after quenching [4]) and are therefore not explicitly included in the present model. The time-dependent evolution of $c^{\text{vac,res}}_{\text{dyn}}$ plays a critical role in determining solute diffusivity and thus strongly influences the kinetics of solute clustering during natural aging. Figs 8b and 8c show the cluster concentration distributions, $c^{\text{clu}}$, as a function of cluster size, $N^{\text{clu}}$, predicted by the CD simulations at various natural aging times: as-quenched (AQ), 1 hour, 6 hours, and 12 hours after the quenching. These results are shown for both fast and slow cooling conditions. Under the fast cooling condition (Fig. 8b), $c^{\text{clu}}(N^{\text{clu}})$ distributions corresponding to small quenched-in clusters are initially present. These clusters continue to grow and coarsen during natural aging due to the elevated diffusivity enabled by higher vacancy supersaturation in the Al matrix. In contrast, the slow cooling condition (Fig. 8c) leads to the formation of larger initial clusters and a substantially reduced concentration of mobile vacancies. As a result, the cluster size distribution evolves much more slowly during natural aging. These results demonstrate that the rate of cluster evolution during aging is strongly influenced by the preceding thermal history, with faster cooling leading to more rapid diffusion and clustering kinetics due to a greater degree of vacancy supersaturation.
Discussions
In this study, we developed an integrated computational framework that combines atomistic KMC simulations with continuum-scale CD modeling, centered on an absorbing Markov chain model, to quantitatively investigate vacancy-cluster interactions in Al-Mg-Zn alloys and their impact on long-time natural aging kinetics. While the absorbing Markov chain approach has been previously applied to address energy basin challenges in computational materials science [19, 20, 21], we specifically adapted it here to explore vacancy trapping in multicomponent alloy systems. Our model accurately predicts vacancy escape times ($t_{\text{esc}}$) from Mg-Zn-rich clusters across a wide range of cluster sizes and configurations. It is parameterized using energetics from surrogate models trained on first-principles calculations (Figs. 2b-c) [54, 30], explicitly incorporating both thermodynamic driving forces via local site binding energies ($E_{\text{bind}}^{\text{site}}$) and kinetic barriers through local bond migration energies ($E_{\text{mig}}^{\text{bond}}$). This approach provides a comprehensive description of the so-called “vacancy prison” effect induced by solute clusters (Fig. 2e). Importantly, we rigorously validated the model against direct KMC simulations (Fig. 4), demonstrating its accuracy and the robustness of the simplified PEL assumption, even in the presence of atomic rearrangements within multicomponent clusters.
The absorbing Markov chain model developed in this work represents a significant advancement in understanding vacancy-cluster interactions, particularly in reconciling previously conflicting descriptions of effective vacancy binding energies. Although the importance of vacancy-cluster binding in solute clustering has been widely recognized, earlier studies [33, 55, 56, 57, 58, 59, 10, 60, 61, 8, 9, 53, 11] have provided contrasting models for how the binding energy evolves with cluster size. For instance, Zurob and Seyedrezai [33] proposed a linear relationship between vacancy binding energy and cluster size, suggesting $E^{\text{clu}}_{\text{bind}}\propto N^{\text{clu}}E^{\text{X}}_{\text{bind}}$, where $N^{\text{clu}}$ is the number of solute atoms in the cluster and $E^{\text{X}}_{\text{bind}}$ is the vacancy binding energy of an individual solute atom. In contrast, Soisson et al. [10] and Vincent et al. [9] studied Cu clusters in $\alpha$-Fe and found that the effective binding energy saturates to a value determined by the difference in vacancy formation energies between the matrix and the precipitate: $E^{\text{clu}}_{\text{bind}}=E_{\text{form}}^{\text{vac,ppt}}-E_{\text{form}}^{\text{vac,mx}}$.
Our results, derived from the absorbing Markov chain model (Fig. 5), reveal a continuous size-dependent transition in the effective vacancy binding energy, $\tilde{E}^{\text{clu}}_{\text{bind}}$, that naturally reconciles these perspectives. For small clusters (up to roughly 100 atoms), $\tilde{E}^{\text{clu}}_{\text{bind}}$ increases sharply with cluster size, reflecting a rapid deepening of the “vacancy prison” energy basin as more solute atoms contribute binding sites. This regime aligns with the linear behavior described by Zurob and Seyedrezai [33]. As the cluster size increases further, the growth of $\tilde{E}^{\text{clu}}_{\text{bind}}$ slows and gradually saturates, consistent with the mechanism proposed by Soisson et al. [10] and Vincent et al. [9], where the binding energy reflects the intrinsic difference in vacancy formation energies between the cluster and the matrix. Importantly, our model also captures the continued, albeit slower, increase in the vacancy escape time, $t_{\text{esc}}$, and thus in $\tilde{E}^{\text{clu}}_{\text{bind}}$, even in the saturation regime. This increase arises from the growing number of random walk steps required for a vacancy to diffuse out from the core of larger clusters. Overall, this continuous transition, inherently captured by the absorbing Markov chain model, demonstrates that previous models effectively described different stages of the same underlying physical phenomenon.
The Markov chain model developed in this study plays a crucial role in enabling mesoscale simulations of early-stage precipitation kinetics. A key challenge in accurately modeling this behavior lies in quantifying the dynamic concentration of vacancies in the presence of solute clusters formed during quenching. This is addressed by incorporating the effective vacancy binding energies (Eq. 13) into the CD framework (Fig. 6). Our predicted vacancy concentrations show good agreement with experimental observations, which consistently report that faster cooling rates lead to higher residual vacancy concentrations [62, 51]. This vacancy concentration correction is essential, as it directly governs solute diffusivity ($D^{\text{X}}$), nucleation rate ($dN/dt$), and growth kinetics ($dr/dt$) during aging [4, 5, 36]. The predictive capability of this approach is validated by excellent agreement with direct KMC simulations (Fig. 7). Furthermore, long-time CD simulations (Fig. 8) reveal that quenching conditions have a lasting impact on the rate of cluster evolution during natural aging. This result provides a quantitative explanation for extensive experimental findings that slow cooling rates degrade mechanical properties such as strength and toughness [63, 64, 65, 66, 67, 68]. The degradation arises because slow cooling promotes the formation of coarse, quench-induced precipitates, which consume both solute atoms and excess vacancies, limiting the formation of fine, strengthening precipitates during subsequent aging [69, 62, 51].
Despite the advances enabled by our integrated multiscale framework, some limitations remain in the current implementation. In particular, our CD simulations, although informed by KMC and Markov chain model results, are performed using a fixed Zn/Mg ratio of 2. This simplification is adopted to avoid the computationally prohibitive task of exploring the full two-dimensional compositional space defined by the Mg and Zn atomic fractions ($n_{\text{Mg}}$, $n_{\text{Zn}}$) in cluster formation enthalpy calculations. The selected ratio is physically meaningful, corresponding to the minimum cluster formation enthalpy pathway (see Supplementary Note 2) and consistent with recent thermodynamic assessments of Mg-Zn clustering in Al alloys [70]. While this approach does not capture the dynamic evolution of the Zn/Mg ratio, which is known to vary with alloy composition and thermomechanical processing conditions [71, 72, 73, 74], it remains a reasonable approximation within the scope of this study. Specifically, our focus is on vacancy trapping effects and their impact on long-term aging kinetics, where this simplification is justified by two key observations. First, at the atomistic level, we find that vacancy trapping behavior is comparable between clusters derived from SA simulations with a fixed Zn/Mg ratio of 2 and those from KMC simulations with varying Zn/Mg ratios (see Supplementary Note 3). Second, at the mesoscale, our simplified CD model accurately reproduces the system energy evolution and cluster size distributions predicted by direct KMC simulations (Fig. 7). These results collectively support the validity of the fixed-ratio approximation for the purposes of this study.
Additionally, while our CD simulations effectively model long-term cluster evolution kinetics, they remain inherently limited by assumptions from classical nucleation theory (CNT) and simplified growth models, such as single-atom addition mechanisms. These simplifications restrict the ability of CD to capture complex phenomena such as coarsening, cluster morphology evolution, and microstructural interactions. To overcome these limitations, larger-scale simulation techniques, such as PF modeling [22, 23, 24], are essential. PF methods are well-suited for simulating microstructural evolution and morphological changes across larger spatial and temporal scales. The predictive capability of PF simulations, however, depends critically on the accurate parameterization of the free energy functional and mobility terms. Our integrated computational framework provides key inputs for such models: surrogate models trained on first-principles data yield reliable thermodynamic quantities for free energy landscapes, while the dynamic vacancy concentrations derived from our approach offer corrections to solute diffusivities and mobilities required in PF formulations. This integration opens pathways for more comprehensive, large-scale simulations of precipitation and microstructure evolution in engineering alloys.
In conclusion, the absorbing Markov chain framework presented here successfully bridges atomistic vacancy interaction mechanisms with macroscopic precipitation behavior. The derived insights into vacancy-cluster interactions and the refined modeling of dynamic vacancy concentrations establish a robust foundation for understanding clustering kinetics, predicting precipitation behaviors, optimizing heat treatments, and ultimately enhancing the performance of engineering alloys such as the Al-Mg-Zn system.
Methods
Kinetic Monte Carlo simulations
The details of the KMC simulation methods are provided in our recent studies and are briefly summarized below [30]. The most common approach to KMC simulations is the first-order residence time algorithm [75]. In this algorithm, we consider $z=12$ first nearest neighbors of a vacancy. Assuming each neighboring atom vibrates at a typical crystal frequency ($\nu_{i}=1\times 10^{13}$ Hz), it results in a transition probability per unit time $r_{i}=\nu_{i}\exp(-E^{\text{bond}}_{\text{mig},i}/k_{B}T)$, with $E^{\text{bond}}_{\text{mig},i}$ representing the migration energy barrier. However, low-barrier events near the solute clusters often lead to forward and backward vacancy jumps, causing frequent ’flicker’ events that considerably slow down first-order KMC algorithms, particularly at low temperatures. We employed a second-order residence time algorithm [38, 39, 41], which bypasses flicker events by considering two-step vacancy jumps. For solute clustering studies, KMC simulations were performed in a $30\times 30\times 30$ FCC supercell of Al-2.86 at.%Mg-2.38 at.%Zn-0.001 at.%vacancy, representative of 7075 Al alloy compositions. The solute concentrations are within the range of compositions of the 7075 Al alloys. Initial configurations were equilibrated at 800 K (typical solutionizing temperature for Al alloys) via CMC for approximately 109 steps. During simulated quenching, temperatures were updated at each timestep according to predefined cooling profiles. For the absorbing Markov chain model verification, simulations employed the same $17\times 17\times 17$ FCC supercell containing clusters obtained from SA and one vacancy, conducted at a higher temperature of 450 K for simulation efficiency.
Absorbing Markov chain
Figure 9: The transition graph of an absorbing Markov chain with 5 states, where states 1, 2, and 3 are transient states and states 4 and 5 are absorbing states.
A Markov chain is a model that describes a system transitioning between a set of states with certain probabilities. It assumes that the next state depends only on the current state, not the past sequence of states. The system is defined by a set of positions (all possible states), a transition matrix (probabilities of moving from one state to another). Monte Carlo (MC) simulations closely relate to Markov chains as both model stochastic system evolution via states with Markovian transition probabilities. The transition matrix $\boldsymbol{P}$ is a matrix of all transition rates $P_{ij}$. Here $P_{ij}$ is the probability for state $i$ to state $j$. The summation of all elements in each row should be 1, we have $\sum_{j}P_{ij}=1$.
An absorbing Markov chain is a type of Markov chain that includes at least one absorbing state, which is defined as a state that, once entered, cannot be exited. This means the probability of remaining in the absorbing state is 1, i.e., $P_{ii}=1$, and the probability of transitioning to any other state is 0, i.e., $P_{ij}=0$ for $i\neq j$. In contrast, a transient state is a state that can be exited, making it possible to leave and potentially never return. The transition matrix $\boldsymbol{P}$ of an absorbing Markov chain can be expressed in canonical form as:
$\boldsymbol{P}=\begin{pmatrix}\boldsymbol{T}&\boldsymbol{R}\\ \boldsymbol{0}&\boldsymbol{I}\end{pmatrix},$ (14)
where $\boldsymbol{I}$ is a $n_{\text{ab}}\times n_{\text{ab}}$ identity matrix corresponding to the absorbing states, with $n_{\text{ab}}$ is the number of absorbing states. $\boldsymbol{T}$ is a $n_{\text{tr}}\times n_{\text{tr}}$ transient matrix containing the probabilities of transitions between transient states. $n_{\text{tr}}$ is the number of transient states. $\boldsymbol{R}$ is a $n_{\text{tr}}\times n_{\text{ab}}$ matrix, representing transition probabilities from transient to absorbing state. $\boldsymbol{0}$ is a zero $n_{\text{ab}}\times n_{\text{tr}}$ matrix indicating no transitions from absorbing to transient states.
Fig. 9 shows an example graph of an absorbing Markov chain with 5 states, where states 1, 2, and 3 are transition states and states 4 and 5 are absorbing states. Eq. 14 for this specific graph is:
$\boldsymbol{P}_{\text{example}}=\begin{pmatrix}0&P_{12}&P_{13}&P_{14}&0\\ P_{21}&0&P_{23}&0&P_{25}\\ P_{31}&P_{32}&0&P_{34}&P_{35}\\ 0&0&0&P_{44}&0\\ 0&0&0&0&P_{55}\end{pmatrix}.$ (15)
Cluster dynamics simulations
CD simulations bridge atomistic simulations and macroscopic approaches by tracking the evolution of clusters in an alloy [44, 45, 46, 47, 48]. In this formalism, clusters with size of $N^{\text{clu}}=n$ atoms were explicitly tracked via their concentration $c_{n}$, governed by a set of coupled ordinary differential equations known as the master equations:
$\frac{dc_{n}}{dt}=(\alpha_{n+1}c_{n+1}+\beta_{n-1}c_{n-1})-(\alpha_{n}+\beta_{n})c_{n},\quad\text{for }n\geq 2,$ (16)
where $\alpha_{n}$ represents the evaporation rate (the rate of cluster shrinkage from size $n$ to $n-1$) and $\beta_{n}$ denotes the condensation rate (the rate of cluster growth from size $n$ to $n+1$). The concentration of monomers (single atoms), $c_{1}$, is constrained by the conservative condition of the total solute concentration $c_{\text{tot}}$:
$c_{1}=c_{\text{tot}}-\sum_{n\geq 2}{nc_{n}},$ (17)
With the quasi-equilibrium assumption at precipitate-matrix interfaces that determines the driving force of diffusion flux [76, 77], these rates present a relationship:
$\frac{\beta_{n}}{\alpha_{n+1}}=c_{1}\exp\left(-\frac{\Delta G_{n+1}-\Delta G_{n}-\Delta G_{1}}{k_{B}T}\right)$ (18)
where $\Delta G_{n}$ is the free energy of formation for a cluster of size $n$, composed of Mg and Zn atoms:
$\Delta G_{n}=\Delta H_{n}-T\Delta S_{n}$ (19)
Although the accurate description of multicomponent clusters in Al-Mg-Zn alloys requires treating $\Delta G_{n}$ as a function of both $n_{\text{Mg}}$ and $n_{\text{Zn}}$, we adopt a fixed Zn/Mg ratio of 2 to reduce complexity and ensure computational efficiency. This Zn/Mg ratio in Mg-Zn clusters is almost consistent with the recent findings of Liu et al. [70]. The enthalpy of formation $\Delta H_{n}$ is modeled using a fitting function [78]:
$\Delta H_{n}=\begin{cases}An-An^{p}&\text{for }n\leq q,\\ Bn+Cn^{1/2}+D&\text{for }n>q,\end{cases}$ (20)
where $A=-0.969$ eV, $B=-0.250$ eV, $C=0.605$ eV, $D=-0.044$ eV, $p=0.952$, and $q=63$ are fitting parameters from the SA simulation results. For the entropy term $\Delta S_{n}$, we only consider the configurational entropy contributions when solute atoms are in dilute solid solutions:
$\Delta S_{n}\approx k_{B}\left[\frac{n}{3}\ln\left(c_{\text{tot}}^{\text{Mg}}\right)+\frac{2n}{3}\ln\left(c_{\text{tot}}^{\text{Zn}}\right)\right]$ (21)
$\beta_{n}$ were explicitly modeled as [79, 80]:
$\beta_{n}=\frac{B_{11}B_{22}-{B_{12}}^{2}}{B_{11}\sin^{2}\theta+B_{22}\cos^{2}\theta},$ (22)
where angle $\theta$ denotes the nucleation and growth direction in the $(n_{\text{Mg}},n_{\text{Zn}})$ space. With Zn/Mg ratio of 2, $\theta=\arctan 2$. $\boldsymbol{B}$ is the matrix characterizing the diffusion-limited condensation process for the critical cluster [45]:
$\boldsymbol{B}=\frac{4\pi R_{n}}{\Omega}\begin{pmatrix}{c_{1}^{\text{Mg}}D^{\text{Mg}}}&0\\ 0&c_{1}^{\text{Zn}}D^{\text{Zn}}\end{pmatrix},$ (23)
where $R_{n}$ is the radius of a solute cluster containing $n$ atoms by assuming it is in a spherical shape, $\Omega$ is the Wigner-Seitz cell volume (the volume of the single atomic site). Both $R_{n}$ and $\Omega$ were calculated based on an FCC lattice with a lattice constant of 4.046Å. $c_{1}^{\text{Mg}}$ and $c_{1}^{\text{Zn}}$ are the concentration of single solute Mg and Zn, respectively. $D^{\text{Mg}}$ and $D^{\text{Zn}}$ are the solute diffusion parameters of Mg and Zn in Al:
$D^{\text{X}}=D^{\text{X}}_{0}\exp\left(-{Q^{X}}/{k_{B}T}\right)\frac{c^{\text{vac,res}}_{\text{dyn}}}{c^{\text{vac,Al}}_{\text{eq}}},$ (24)
Wu et al. [81] provided first-principles solute diffusion constants, $D^{\text{X}}_{0}$ for Mg of $9.4182\times 10^{-2}$ cm2/s and Zn of $1.0614\times 10^{-1}$ cm2/s, and solute diffusion activation energy, $Q^{X}$, for Mg and Zn in Al matrix of $1.2467$ eV and $1.1812$ eV, respectively. $c^{\text{vac,res}}_{\text{dyn}}$ is calculated based on Eq. 13 using parameters $\tilde{E}_{\text{bind}}^{\text{clu}}$ obtained from KMC and absorbing Markov chain models. $c^{\text{vac,Al}}_{\text{eq}}$ is the equilibrium vacancy concentration:
$c^{\text{vac,Al}}_{\text{eq}}=c_{0}\exp(\frac{S^{\text{vac,Al}}}{k_{B}})\exp(-\frac{H^{\text{vac,Al}}}{k_{B}T})$ (25)
where we adopt the values from pure Al [3]: $c_{0}=1.67$ as the preexponential factor, $H^{\text{vac,Al}}=0.66$ eV as the vacancy formation enthalpy and $S^{\text{vac,Al}}=0.7k_{B}$ as the vacancy formation entropy.
Acknowledgements
This research was supported by NSF grant #1905421, with computing resources provided by GM Global Technical Center and Stampede3 at TACC (ACCESS allocation DMR190035, NSF grants #2138259, #2138286, #2138307, #2137603, #2138296). The authors thank insightful discussions with Prof. Dane Morgan from the University of Wisconsin.
Author contributions statement
Z.X. and L.H. performed the high-throughput first-principles simulations. Z.X. and L.Q. constructed the models and analyzed the results. L.H., A.M., and L.Q. conceptualized and supervised the research project. All authors analyzed the results and prepared the manuscript together.
Competing interests
The authors declare no competing interests.